OpenStudy (anonymous):

Could anyone please help me with simplifying radicals? -15

OpenStudy (anonymous):

√(-15) is the problem?

OpenStudy (anonymous):

\[\sqrt{-15}\]

OpenStudy (anonymous):

what is your question?

OpenStudy (anonymous):

it is odd and cannot be simplfied

OpenStudy (anonymous):

ok so if you break it into it's factors you get √(-1)(3)(5) √(-1) = i the rest cannot be simplified so you have i√15 for an answer

OpenStudy (anonymous):

√49 is also odd but simplifies to 7 mmolina

OpenStudy (anonymous):

so the answer is -1/15

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no, the answer is: i*srqr(15)

OpenStudy (anonymous):

i squared as in india type of i?

OpenStudy (anonymous):

\[i*\sqrt{15}\]

OpenStudy (anonymous):

hmm never heard of that before. what rule is that derived from?

OpenStudy (anonymous):

i is the imaginary unit √(-1) because you cannot take the square root of a negative number.

OpenStudy (anonymous):

okay. let me try that...

OpenStudy (anonymous):

set in you personal calculator \[\sqrt{-1}\] this will give you an error as the answer. So every time you have an expresion like that you have to write an i in front of the root and make the number inside the root positive.

OpenStudy (anonymous):

the entire problem is quadratic formulas which is -1 +/- \[\sqrt{-15}/2\]

OpenStudy (anonymous):

then there are no real solutions

OpenStudy (anonymous):

I have the left side down and the fraction beneathe, but complicated fixing right side. is that possible for a solution set? the answer must be two separate numbers

OpenStudy (anonymous):

not always, quadratics have AT MOST 2 solutions but can have only one or no real solutions, this appears to be a classic no real solutions problem

OpenStudy (anonymous):

the solutions to a quadratic are where the graph crossed the x-axis, if you graph the original problem does it cross the x-axis?

OpenStudy (anonymous):

there is no graphing, and there is no option for no solution--there is indeed a solution. haha

OpenStudy (anonymous):

It might make things simpler if you tell us what the original problem is.

OpenStudy (anonymous):

it must form an equation with only addition and subraction signs between them

OpenStudy (anonymous):

I did if you look above

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

did you not understand it?

OpenStudy (anonymous):

I understand that you have simplified the original problem but if you post what the quadratic is then we could verify that your solutions are correct.